# On the Resilience of LLM-Based Multi-Agent Collaboration with Faulty Agents

### v3

@@ Abstract
Multi-Agent System, Large Language Model, Resilience, Robustness
@@ 1 Introduction
Multi-agent collaboration has further boosted Large Language Models’ (LLMs) already impressive performance across various downstream tasks, including code generation (Liu et al., 2023; Lee et al., 2024), math problem solving (Lu et al., 2024; Liang et al., 2024), and text translation (Jiao et al., 2023; Wu et al., 2024a). Multi-agent systems achieve these improvements by decomposing complex tasks into smaller, specialized sub-tasks handled by expert agents (Chen et al., 2024b; Li et al., 2024).
@@
However, the decentralized nature of multi-agent systems leaves them vulnerable to clumsy or malicious agents, which could undermine or destroy collaboration (Chen et al., 205). Consider a scenario where companies specializing in different areas produce expert agents, the lack of centralized control means that the multi-agent system may contain agents from various sources, some of which could be faulty. In a multi-agent coding system like Camel (Li et al., 2023), a faulty coding agent produces buggy code, causing severe errors or harmful outputs when executed by other agents.
Recent studies (Zhang et al., 2024; Tian et al., 2023; Amayuelas et al., 2024; Ju et al., 2024; Yu et al., 2024) have increasingly focused on safety issues within multi-agent systems. However, these studies mainly investigate attacks on agents to induce toxicity in their outputs or misinformation spread among all agents. While they assess malicious agent behavior against safety benchmarks like AdvBench (Zou et al., 2023), they overlook the disruption of collaboration in solving general tasks and the impact of varying system structures. In this paper, we study the resilience of multi-agent collaboration against faulty agents, specifically the systems’ ability to recover from errors.
First, we propose two approaches to simulating agents’ faulty behaviors on various tasks, namely AutoTransform and AutoInject. AutoTransform transforms a given agent’s profile into a faulty version that retains original functionalities while introducing stealthy errors. AutoInject is designed to directly and automatically inject errors into messages spread among agents. The two methods offer automate introduction of errors in multi-agent collaboration without requiring manual modifications.
@@
Our key findings include: (1) The Hierarchical structure exhibits the least performance drop at $9.2\%$, aligning with its prevalence in human organizational structures (Mihm et al., 2010). (2) Code Generation, as a relatively objective task, is most affected by malicious agents, experiencing a performance drop of $24.7\%$. (3) Manually introducing errors can sometimes improve the overall performance, especially on MAD (Liang et al., 2024). (4) Increasing the ratio of Faulty Messages and using Semantic Errors results in a greater performance drop than increasing the number of errors per message and using syntactic errors. (5) The combination of The Challenger and The Inspector enhances system resilience most for the two more vulnerable systems: Self-collab with a linear structure and Camel with a flat structure, recovering up to $96.4\%$ of performance lost caused by faulty agents. The contribution of this paper are as follows:
@@ 4.1 Experimental Settings
Not all systems are designed to support the four tasks studied in our paper. Therefore, we modified the prompts of some systems to adapt to our selected tasks. The modified prompts are detailed in §C.1. GPT-3.5 is consistently used for both AutoTransform and AutoInject to ensure a fair comparison. We use GPT-3.5 and GPT-4o as the backbone for these systems for main experiments (RQ1 and RQ2) while using GPT-3.5 for factor analysis. All LLMs are used with a temperature of zero. We introduce one faulty agent at a time to avoid interference and facilitate essential analysis, which is shown in Table 1. Normal agents remain unaware of the faulty agent’s presence, reflecting a realistic information-asymmetric scenario (Zhou et al., 2024a).
@@ 4.2 RQ1: Impact of System Structures
[…]
AutoInject causes a larger performance drop than AutoTransform on GPT-3.5, but a lower performance drop using GPT-4o. While one might assume AutoTransform would have a greater negative impact on multi-agent collaboration due to its permanent modification of agents’ profiles into faulty ones, it is AutoInject that results in a larger performance drop using GPT-3.5. The reasons for this are two-fold: (1) Current LLMs have a weakness where they become less effective as the context lengthens, especially where conflict exists in instructions. For our faulty agents, they gradually lose track of the task to produce errors, prioritizing new instructions from other agents to correct errors in the message. (2) AutoInject consistently introduces errors, whereas AutoTransform does not always ensure error generation. Despite being transformed into faulty agents, they sometimes fail to generate errors due to constraints requiring errors to be stealthy. These issues are mitigated as the capabilities of LLMs advance. With GPT-4o as the backbone, the faulty agents generated by AutoTransform demonstrate a strong capacity for instruction following, resulting in stealth errors that lead to a more significant performance decline compared to AutoInject.
@@
Figure 4: The performance of various tasks with the two error-introducing methods, with results averaged across all six multi-agent systems (three system structures).
4.3 RQ2: Impact of Downstream Tasks
Tasks requiring rigor and formalization, such as code generation and math, are more sensitive to agent errors and exhibit lower resilience compared to translation and text evaluation. Code generation and math demand greater objectivity than the more subjective tasks of translation and text evaluation. Fig. 4a and 4b illustrate the impact of AutoTransform and AutoInject on different downstream tasks, averaged across all multi-agent systems. We also present the performance of single-agent using the prompts listed in §C.2, for a clearer comparison. The results indicate several conclusions: (1) Multi-agent systems can outperform single-agent settings, but their performance may decline to similar or worse levels when affected by faulty agents. (2) Objective tasks benefit more from multi-agent collaboration, while subjective tasks gain less. Additionally, errors in subjective tasks are often overlooked by other agents due to the lack of rigorous correctness standards. (3) In terms of system resilience, tasks ranked from least to most vulnerable are: code generation, math, translation, and text evaluation. Even minor errors in the first two tasks, particularly in code generation, greatly affect rigor and formalization. Conversely, the latter two tasks are less sensitive to minor variations in a single agent’s output. (4) AutoTransform decreases more performance than AutoInject except in code generation using GPT-3.5.
@@ 4.3 RQ2: Impact of Downstream Tasks
(a) Using different error rates with either $P_{e}$ or $P_{m}$ fixed.
(b) Using either semantic or syntactic errors.
4.4 RQ3: Impact of Error Rates
@@ 6.1 Multi-Agent Systems
LLMs enhance multi-agent systems through their exceptional capability for role-play (Wang et al., 2024b). Despite utilizing a same architecture like GPT-3.5, tasks benefit from tailored in-context role-playing prompts (Min et al., 2022). Besides the six frameworks selected in this study, researchers have been exploring multi-agent collaboration in downstream tasks or simulated communities. ChatEval (Chan et al., 2024) is a multi-agent debate system for evaluating LLM-generated text, providing a human-like evaluation process. ChatDev (Qian et al., 2024) uses a linear structure of several roles to address code generation tasks. AutoGen (Wu et al., 2024b) offers a generic framework for building diverse applications with multiple LLM agents. AutoAgents (Chen et al., 2024a) enables dynamic generation of agents’ profiles and cooperation, evaluated on open-ended QA and creative writing tasks. Zhou et al. (2023) support planning, memory, tool usage, multi-agent communication, and fine-grained symbolic control for multi-agent or human-agent collaboration. Additionally, there are studies simulating daily life or conversations (Park et al., 2023; Zhou et al., 2024b; Yang et al., 2024), multi-agent competition (Huang et al., 2025; Liu et al., 2024; Liang et al., 2023), or agentic workflow (Qian et al., 2025; Zhuge et al., 2024). These frameworks are not selected either because they are not task-oriented (e.g., simulated society or competitions) or their system design overlaps with those chosen for this study.
@@ 6.2 Safety Issues in Multi-Agent Systems
PsySafe (Zhang et al., 2024) is a framework that integrates attack, evaluation, and defense mechanisms using psychological manipulation involving negative personalities. EG (Evil Geniuses) (Tian et al., 2023) is an attack method that automatically generates prompts related to agents’ original roles, similar to our AutoTransform. While PsySafe and EG are applied to different multi-agent systems such as Camel and MetaGPT, they do not examine the impact of adversaries on downstream tasks like code generation or translation. Agent Smith (Gu et al., 2024) showed that malicious behaviors can spread among agents, using multi-agent interaction and memory storage. Similarly, Yu et al. (2024) used adversarial attack to jailbreak all agents with a single message from a single agent. Amayuelas et al. (2024) investigates how an adversary in multi-agent debate can disrupt collaboration in tasks including MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), MedMCQA (Pal et al., 2022), and LegalBench (Guha et al., 2023), finding that the adversary’s persuasion skill is crucial for a successful attack. Ju et al. (2024) proposes a two-stage attack strategy to create an adversary that spreads counterfactual and toxic knowledge in a simulated multi-agent chat environment. This method can effectively break collaboration in MMLU. Unlike our study, Amayuelas et al. (2024) and Ju et al. (2024) do not explore how different system architectures are affected by these adversaries.
@@ Acknowledgments
The paper is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490410. The paper is also supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14206921 of the General Research Fund).
Appendix A Error Type Analysis
Category Name Description Count
Appendix B Quantitative Results
@@
Task Linear Flat Hierarchical
GPT-3.5 55.62 54.37 53.00
        w/ AutoTransform 38.24 43.93 46.57
        w/ AutoInject 38.27 40.25 48.12
GPT-4o 67.18 67.52 67.79
        w/ AutoTransform 30.08 56.92 60.83
        w/ AutoInject 44.14 60.51 64.01
@@
Task Code Gen Math Translation Text Eval
GPT-3.5 64.70 30.00 69.10 45.45
        w/ AutoTransform 50.83 22.67 67.99 43.68
        w/ AutoInject 39.1 25.00 67.85 46.50
GPT-4o 81.91 60.00 70.82 54.17
        w/ AutoTransform 73.78 50.67 65.12 52.92
        w/ AutoInject 72.22 51.33 71.36 54.33
@@
Model MetaGPT Self-collab Camel SPP MAD AgentVerse Average
@@
Model MetaGPT Self-collab Camel SPP MAD AgentVerse Average

### v4

@@ Abstract
Multi-Agent Systems, Large Language Models, Resilience, Robustness
@@ 1 Introduction
Multi-agent systems have further boosted Large Language Models’ (LLMs) already impressive performance across various downstream tasks, including code generation (Liu et al., 2023; Lee et al., 2024), math problem solving (Lu et al., 2024; Liang et al., 2024), and text translation (Jiao et al., 2023; Wu et al., 2024a), by decomposing complex tasks into smaller, specialized sub-tasks handled by expert agents (Chen et al., 2024b; Li et al., 2024). However, the decentralized nature of multi-agent systems leaves them vulnerable to clumsy or malicious agents, which could undermine or destroy collaboration (Chen et al., 2025b). Consider a scenario where companies specializing in different areas produce expert agents, the lack of centralized control means that the multi-agent system may contain agents from various sources, some of which could be faulty. In a multi-agent coding system like Camel (Li et al., 2023), a faulty coding agent produces buggy code, causing severe errors or harmful outputs when executed by other agents.
Recent studies (Zhang et al., 2024; Tian et al., 2023; Amayuelas et al., 2024; Ju et al., 2024; Yu et al., 2024c) have increasingly focused on safety issues within multi-agent systems. However, these studies mainly investigate attacks on agents to induce toxicity in their outputs or misinformation spread among all agents. While they assess malicious agent behavior against safety benchmarks like AdvBench (Zou et al., 2023), they overlook the disruption of collaboration in solving general tasks and the impact of varying system structures on overall resilience.
In this paper, we study the resilience of multi-agent collaboration against faulty agents, specifically the systems’ ability to recover from errors. First, to simulate agents’ faulty behaviors across various tasks with precise control over error rates and types, we propose two approaches: (1) AutoTransform transforms a given agent’s profile into a faulty version that retains original functionalities while introducing stealthy errors. (2) AutoInject is designed to directly and automatically inject errors into messages spread among agents. The two methods offer automate introduction of errors in multi-agent collaboration without requiring manual modifications.
@@
Our key findings include: (1) The Hierarchical structure exhibits the least performance drop at $5.5\%$, aligning with its prevalence in human organizational structures (Mihm et al., 2010). (2) Code Generation, as a relatively objective task, is most affected by malicious agents, experiencing a performance drop of $22.6\%$. (3) Manually introducing errors can sometimes improve the overall performance, especially on MAD (Liang et al., 2024). (4) Increasing the ratio of Faulty Messages and using Semantic Errors results in a greater performance drop than increasing the number of errors per message and using syntactic errors. (5) The combination of The Challenger and The Inspector enhances system resilience most for the two more vulnerable systems: Self-collab with a linear structure and Camel with a flat structure, recovering up to $96.4\%$ of performance lost caused by faulty agents. The contribution of this paper are as follows:
@@ 2 Preliminaries
Formally, we can a multi-agent system as a graph: $G=(V,E)$, where $V$ represents agents and $E\subseteq V\times V$ is a set of directed edges. Each $(u,v)\in E$ denotes agent $u$ reports to agent $v$. (1) Linear systems are directed path graphs, where $\forall v\in V,v\neq s,v\neq t$, we have: $\deg^{+}(v)=\deg^{-}(v)=1$; for the endpoints, $\deg^{-}(s)=0$, $\deg^{+}(s)=1$, $\deg^{+}(t)=0$, and $\deg^{-}(t)=1$. Agents in this structure form a chain from $s$ to $t$. (2) Flat systems are directed complete graphs with bidirectional edges, where $\forall u,v\in V,u\neq v$, both $(u,v)\in E$ and $(v,u)\in E$. This represents a fully connected, non-hierarchical structure. (3) Hierarchical systems are rooted directed trees, where there exists a unique root agent $r\in V$ such that $\deg^{-}(r)=0$, and $\forall v\in V\setminus{r}$, $\deg^{-}(v)=1$. The structure is acyclic and forms a strict top-down hierarchy.
@@ 4 Experiments
Not all systems are designed to support the four tasks studied in our paper. Therefore, we modified the prompts of some systems to adapt to our selected tasks. The modified prompts are detailed in §C.1. GPT-3.5 is consistently used for both AutoTransform and AutoInject to ensure a fair comparison. We use GPT-3.5 and GPT-4o as the backbone for these systems for main experiments (RQ1 and RQ2) while using GPT-3.5 for factor analysis. All LLMs are used with a temperature of zero. We introduce one faulty agent at a time to avoid interference and facilitate essential analysis, which is shown in Table 1. Normal agents remain unaware of the faulty agent’s presence, reflecting a realistic information-asymmetric scenario (Zhou et al., 2024a). All the systems adopt direct messaging as the communication scheme and no other message processing schemes such as summarizing or broadcasting is involved.
@@ 4.1 RQ1: Impact of System Structures
The hierarchical structure has a higher resilience than other two, exhibiting the smallest accuracy drop. Fig. 3a and 3b illustrate the impact of AutoTransform and AutoInject on various structures of multi-agent system, averaged across different downstream tasks. The ranking of system resilience from strongest to weakest—hierarchical, flat, and linear—is consistent across both GPT-3.5 and GPT-4o, as well as under both error-introducing methods. We attribute this resilience to the presence of a higher-level agent (e.g., the evaluator in MAD), which is always presented with various versions of the answer by multiple agents performing the same sub-task, increasing the likelihood of error recovery from a single agent. The flat structure shows a lower resilience ($-10.54$) than the hierarchical structure ($-5.51$), while the linear architecture demonstrates the lowest resilience ($-23.72$). A hierarchical structure enables centralized decision-making, where a top-level role gathers information and efficiently distributes decisions through clear chains of command. In contrast, a flat structure often lacks clear leadership, leading to decision paralysis and coordination issues. A linear structure has a chain of command, but communication is slower, and top leaders have limited oversight of lower levels.
AutoInject causes a larger performance drop than AutoTransform on GPT-3.5, but a lower performance drop using GPT-4o. While one might assume AutoTransform would have a greater negative impact on multi-agent collaboration due to its permanent modification of agents’ profiles into faulty ones, it is AutoInject that results in a larger performance drop using GPT-3.5 ($-12.12$), compared to AutoTransform ($-11.42$). The reasons for this are two-fold: (1) Current LLMs have a weakness where they become less effective as the context lengthens, especially where conflict exists in instructions. For our faulty agents, they gradually lose track of the task to produce errors, prioritizing new instructions from other agents to correct errors in the message. (2) AutoInject consistently introduces errors, whereas AutoTransform does not always ensure error generation. Despite being transformed into faulty agents, they sometimes fail to generate errors due to constraints requiring errors to be stealthy. These issues are mitigated as the capabilities of LLMs advance. With GPT-4o as the backbone, the faulty agents generated by AutoTransform demonstrate a strong capacity for instruction following, resulting in stealth errors that lead to a more significant performance decline ($-18.22$) compared to AutoInject ($-11.28$).
4.2 RQ2: Impact of Downstream Tasks
Tasks requiring rigor and formalization, such as code generation and math, are more sensitive to agent errors and exhibit lower resilience compared to translation and text evaluation. Code generation and math demand greater objectivity than the more subjective tasks of translation and text evaluation. Fig. 4a and 4b illustrate the impact of AutoTransform and AutoInject on different downstream tasks, averaged across all multi-agent systems. We also present the performance of single-agent using the prompts listed in §C.2, for a clearer comparison. The results indicate several conclusions: (1) Multi-agent systems can outperform single-agent settings ($+4.76$ for GPT-3.5 and $+5.29$ for GPT-4o), but their performance may decline to similar or worse levels when affected by faulty agents. (2) Objective tasks benefit more from multi-agent collaboration, while subjective tasks gain less. Additionally, errors in subjective tasks are often overlooked by other agents due to the lack of rigorous correctness standards. (3) In terms of system resilience, tasks ranked from least to most vulnerable are: code generation ($-22.56$), math ($-9.89$), text evaluation ($-5.42$), and translation ($-4.70$). Even minor errors in the first two tasks, particularly in code generation, greatly affect rigor and formalization. Conversely, the latter two tasks are less sensitive to minor variations in a single agent’s output. (4) AutoTransform decreases more performance than AutoInject ($-12.45$ compared to $-8.83$), except in code generation using GPT-3.5.
@@ 4.2 RQ2: Impact of Downstream Tasks
Figure 4: The performance of various tasks with the two error-introducing methods, with results averaged across three system structures (all six multi-agent systems).
(a) Using different error rates with either $P_{e}$ or $P_{m}$ fixed.
(b) Using either semantic or syntactic errors.
4.3 RQ3: Impact of Error Rates
@@ 6.1 Multi-Agent Systems
LLMs enhance multi-agent systems through their exceptional capability for role-play (Wang et al., 2024b). Despite utilizing a same architecture like GPT-3.5, tasks benefit from tailored in-context role-playing prompts (Min et al., 2022). Besides the six frameworks selected in this study, researchers have been exploring multi-agent collaboration in downstream tasks or simulated communities (Tran et al., 2025; Zhang et al., 2025b, a). ChatEval (Chan et al., 2024) is a multi-agent debate system for evaluating LLM-generated text, providing a human-like evaluation process. ChatDev (Qian et al., 2024) uses a linear structure of several roles to address code generation tasks. AutoGen (Wu et al., 2024b) offers a generic framework for building diverse applications with multiple LLM agents. AutoAgents (Chen et al., 2024a) enables dynamic generation of agents’ profiles and cooperation, evaluated on open-ended QA and creative writing tasks. Zhou et al. (2023) support planning, memory, tool usage, multi-agent communication, and fine-grained symbolic control for multi-agent or human-agent collaboration. Additionally, there are studies simulating daily life or conversations (Park et al., 2023; Zhou et al., 2024b; Yang et al., 2024), multi-agent competition (Huang et al., 2025; Liu et al., 2024; Liang et al., 2023), or agentic workflow (Qian et al., 2025b; Zhuge et al., 2024). These frameworks are not selected either because they are not task-oriented (e.g., simulated society or competitions) or their system design overlaps with those chosen for this study.
@@ 6.2 Safety Issues in Multi-Agent Systems
Researchers have moved attention towards the reliability in single (Yu et al., 2024a, 2025; Perez et al., 2022; Tan et al., 2021; Wang et al., 2025a; Chen et al., 2025a) and multi-agent systems (Mao et al., 2025; Zhou et al., 2025; Wang et al., 2025b). PsySafe (Zhang et al., 2024) is a framework that integrates attack, evaluation, and defense mechanisms using psychological manipulation involving negative personalities. EG (Evil Geniuses) (Tian et al., 2023) is an attack method that automatically generates prompts related to agents’ original roles, similar to our AutoTransform. While PsySafe and EG are applied to different multi-agent systems such as Camel and MetaGPT, they do not examine the impact of adversaries on downstream tasks like code generation or translation. Agent Smith (Gu et al., 2024) showed that malicious behaviors can spread among agents, using multi-agent interaction and memory storage. Similarly, Yu et al. (2024c) used adversarial attack to jailbreak all agents with a single message from a single agent. Amayuelas et al. (2024) investigates how an adversary in multi-agent debate can disrupt collaboration in tasks including MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), MedMCQA (Pal et al., 2022), and LegalBench (Guha et al., 2023), finding that the adversary’s persuasion skill is crucial for a successful attack. Ju et al. (2024) proposes a two-stage attack strategy to create an adversary that spreads counterfactual and toxic knowledge in a simulated multi-agent chat environment. This method can effectively break collaboration in MMLU. Unlike our study, these studies do not explore how different system architectures are affected by these adversaries. While NetSafe (Yu et al., 2024b) investigates the impact of different numbers and structures of faulty agents, they do not investigate more subjective tasks like translation and text evaluation. Additionally, we include code generation task, enabling us to study the impact of error types.
@@ Acknowledgments
The paper is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490410. The paper is also supported by the Research Grants Council of the Hong Kong Special Administrative Region, China for Theme-based Research Scheme Project (RGC Ref. No. T43-513/23-N).
Appendix A Quantitative Results
@@
Task Linear Flat Hierarchical Avg.
GPT-3.5 55.62 54.37 53.00 54.33
        w/ AutoTransform 38.24 43.93 46.57 42.91
        w/ AutoInject 38.27 40.25 48.12 42.21
GPT-4o 67.18 67.52 67.79 67.50
        w/ AutoTransform 30.08 56.92 60.83 49.28
        w/ AutoInject 44.14 60.51 64.01 56.22
Avg. $\downarrow$23.72 $\downarrow$10.54 $\downarrow$5.51 $\downarrow$13.26
@@
Task Code Gen Math Translation Text Eval Avg.
GPT-3.5 Single Agent 58.41 24.00 68.42 41.25 48.02
[…]
        w/ AutoTransform 44.85 19.53 65.98 43.08 43.36
        w/ AutoInject 39.15 25.25 67.74 42.69 43.71
GPT-4o Single Agent 78.83 44.00 70.38 48.75 60.49
[…]
        w/ AutoTransform 58.49 39.92 58.69 44.06 50.29
        w/ AutoInject 60.15 44.59 71.09 52.94 57.19
Avg. $\downarrow$22.56 $\downarrow$9.89 $\downarrow$4.70 $\downarrow$5.42 $\downarrow$10.64
@@
Model MetaGPT Self-collab Camel SPP MAD AgentVerse Avg.
@@
Model MetaGPT Self-collab Camel SPP MAD AgentVerse Avg.
@@
Appendix B Additional Results
B.1 Error Type Analysis
Category Name Description Count
While both mechanisms are conceptually complementary, we restrict the quantitative error-type analysis that follows to AutoInject. In practice, GPT-3.5 follows rate-based instructions only very loosely when we use AutoTransform to ask an agent to “inject syntax errors in Xof its code lines.” Targeting 20% and 40% line-level error rates, the realized proportions fluctuate wildly—averaging 1.56 (SD = 3.65, min = 0.00, max = 14.30) and 9.49 (SD = 26.70, min = 0.00, max = 90.10) respectively. The large standard deviations and extreme maxima show that faulty agents created by AutoTransform rarely attain the desired granularity, rendering controlled experiments and fair comparisons impossible. AutoInject, by contrast, allows deterministic control over both the location and amount of corruption, providing the reliable and reproducible error patterns required for the analyses that follow.
B.2 Faulty Single Agent Systems
Table 9: Influence of AutoTransform and AutoInject on single agents.
Task Code Gen Math Translation Text Eval Avg.
GPT-3.5 Single Agent 58.41 24.00 68.42 41.25 48.02
        w/ AutoTransform 3.92 8.00 68.42 18.75 21.66
        w/ AutoInject 15.24 18.00 61.08 32.50 31.71
Table 10: Comparison of the performance of single agent and multi-agent systems.
Structure Single Agent Linear Flat Hierarchical
GPT-3.5 48.02 55.62 54.37 53.00
        w/ AutoTransform 21.66 38.24 43.93 46.57
        w/ AutoInject 31.71 38.27 40.25 48.12
We conduct experiments on applying the two error-introducing methods on a single agent based on GPT-3.5 across all four tasks. The performance is shown in Table 9. Compared to the performance of other multi-agent systems in Table 10, we conclude that all three types of systems have better resilience against both methods compared to a single agent. This is because the systems have other “good” agents for reviewing and testing, identifying the errors made by the faulty agent.
B.3 Multiple Faulty Agents
High-level planners dominate the failure cascade in multi-fault settings. Extending the Math experiments in AgentVerse to two simultaneous faulty agents reveals that compounding errors do not affect all roles equally. Using AutoInject, with only the Solver corrupted, accuracy falls from $28.0$ to $20.0$. However, when the additional faulty agent is the Critic, performance drops further to $14.0$, and when the faulty agent is the Planner, it decreases to just $12.0$. The steep drop corroborates the Planner’s higher influence: because it dictates the team’s global search direction, a single mis-step in its high-level plan propagates irrecoverably through subsequent reasoning, even when other specialists remain normal. Similarly, AutoTransform can also cause performance drop on the system ($16.0$ with one faulty agent to $14.0$ with Solver + Critic, and only $2.0$ with Solver + Planner).
B.4 Advanced Multi-Agent Systems
Star-topology graphs preserve the hierarchy advantage. To verify that our conclusions transfer to richer communication patterns, we implement two four-agent graph frameworks inspired by GPTSwarm (Zhuge et al., 2024) and MacNet (Qian et al., 2025a): a complete graph in which every agent exchanges answers with all peers, and a star graph in which a single leader coordinates three workers. On the Math task with GPT-3.5, the star graph attains $36.0$ in the vanilla setting and retains $30.0$ and $28.0$ under AutoTransform and AutoInject, whereas the complete graph lags behind at $28.0$ to $20.0$ and $16.0$. The single leader once again mitigates error propagation, while the fully-flat peer discussion in the complete graph amplifies faults. These observations confirm that our resilience analysis is applicable to diverse, graph-based frameworks, and they reinforce the central insight that even modest hierarchical oversight noticeably boosts robustness.
B.5 More Realistic and Complex Tasks
Multi-agent collaboration produces executable real-world software—bugs appear when including faulty agents. To probe tasks that demand end-to-end software assembly rather than isolated snippets, we ask agents to build a complete Snake game in pygame. A single GPT-3.5 agent delivers only a shell: the window opens but the snake could not be steered. In contrast, Camel’s two-agent loop generates a fully playable game with food spawning, score keeping, and self-collision logic. When we convert the Assistant into a faulty agent with AutoTransform, the program still launches, yet subtle semantic faults surface—the snake advances at an uncontrollable speed and the arrow-key mapping drifted (e.g., Left triggered an upward move), mirroring the “stealth-but-harmful” failures. Under AutoInject, the injected syntactic glitches (missing pygame.init() and a mismatched surface size) render the game un-executable, again confirming that direct message corruption is more destructive than profile transformation. Importantly, despite these challenges, the multi-agent system retains a playable version in two of the three conditions, whereas the single-agent baseline never produces a controllable game—echoing the performance gap between single versus multi-agent systems.
B.6 More Models
The resilience trend holds for non-GPT backbones and Chain-of-Thought (CoT) prompting. Replacing the GPT series with LLaMA-3.1-70B-Instruct leaves the hierarchy-first ordering intact: under AutoTransform the hierarchical system loses only $9.2$ ($76.15$ to $66.96$), while the flat and linear structures plunge by $37.8$ and $61.9$ respectively. AutoInject paints the same picture, with drops of $20.5$ (hierarchical), $40.2$ (flat), and $35.1$ (linear). A similar pattern emerges when we apply the budget-friendly reasoning model, o1-Mini, on the Math task: hierarchy yields a modest $18$ hit under AutoTransform and just $4$ under AutoInject, whereas linear falls by as much as $64$. Crucially, faulty agents still exact a sizable toll, especially on non-hierarchical topologies. These findings confirm that the structural advantage of a hierarchy is model-agnostic, and that while CoT improves raw performance, it does not in itself immunize multi-agent systems against faulty peers.